# MeshCone: Second-Order Cone Programming for Geometrically-Constrained Mesh Enhancement

### v2

@@ 4 Experiments
The approach is specifically evaluated on 53 meshes from the ShapeNet dataset [chang2015shapenetinformationrich3dmodel], with each mesh representing a distinct object category. This diverse evaluation set spans the full range of ShapeNet taxonomy, including vehicles, furniture, electronics, kitchenware, and many other categories. This setup allows us to assess whether the optimization process can generalize across vastly different structures. In addition, we tested our solution on three meshes generated from the work of [Binninger_2025], which are highly detailed and contain thousands of faces and vertices, making them ideal for a real-world test [Laric2012ThreeDScans]. We sampled 10000 points from these representations to obtain the respective centroids used for alignment between the deformed vertices and the ground truth mesh.
@@ 5 Results
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison on three meshes across GroundTruth (left), TetWeave (middle), and MeshCone (right).
Overall, our experiments demonstrate that the proposed convex optimization framework effectively preserves the intrinsic geometric properties of the input meshes while refining their structural consistency. By relying solely on convex solvers and linear constraints, the method achieves stable and interpretable optimizations without the need for complex non-linear regularization or data-driven priors.
@@ 6.1 Limitations
Future work can reduce these failure modes by combining improved training strategies with some constraints that would refine those areas that weren’t completely solved by the convex optimization method. A post-refinement approach could probably solve this issue.
@@ 8 Conclusions
This work introduced a convex optimization framework for mesh improvement that relies exclusively on linear constraints and mathematically interpretable solvers. Through this research, we demonstrated that high-quality developments can be achieved while preserving the intrinsic geometric and topological properties of the input meshes. Our tests on ShapeNet and ThreeDScans confirmed that the proposal really enhances meshes by improving the connectivity among the vertices and edges while including smoothness into the final representation.

### v3

@@ 4 Experiments
The approach is specifically evaluated on 53 meshes from the ShapeNet dataset [chang2015shapenetinformationrich3dmodel], with each mesh representing a distinct object category. This diverse evaluation set spans the full range of ShapeNet taxonomy, including vehicles, furniture, electronics, kitchenware, and many other categories. This setup allows us to assess whether the optimization process can generalize across vastly different structures. In addition, we tested our solution against Laplacian smoothing on three meshes generated from the work of [Binninger_2025], which are highly detailed and contain thousands of faces and vertices, making them ideal for a real-world test [Laric2012ThreeDScans]. We sampled 10000 points from these representations to obtain the respective centroids used for alignment between the deformed vertices and the ground truth mesh.
@@ 5 Results
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Quality comparison on three meshes: Ground Truth (left), TetWeave (middle left), Laplacian-smoothed TetWeave (middle right), and MeshCone (right).
Convex optimization provides the mathematical foundation for precise mesh refinement, guaranteeing global solutions that preserve geometric integrity—unlike Laplacian smoothing which destroys essential details through over-smoothing. Our framework demonstrates that convex solvers with linear constraints achieve optimal shape preservation and detail retention, offering stable, interpretable alternatives to complex non-convex approaches.
@@ 6.1 Limitations
While the convex optimization framework provides strong theoretical guarantees, it occasionally struggles with preserving extremely high-frequency details in complex geometric regions. These limitations could be addressed in future work through specialized constraint formulations that explicitly target fine-scale features.
@@ 8 Conclusions
This work introduced a convex optimization framework for mesh improvement that relies exclusively on linear constraints and mathematically interpretable solvers. Through this research, we demonstrated that high-quality developments can be achieved while preserving the intrinsic geometric and topological properties of the input meshes. Our tests on ShapeNet and ThreeDScans confirmed that the proposal really enhances meshes by improving the connectivity among the vertices and edges while including smoothness into the final representation. MeshCone enables user to refine meshes without requiring complex hyperparameters or long training times, what makes it suitable for different applications related to geometry processing and meshing.